# Bootstrapping Estimators based on the Block Maxima Method

### v1

@@ Appendix B Runtime comparison
The above heuristic has been confirmed by Monte Carlo simulations. Exemplary results are presented in Figure 11, which rely on simulated data from Model 6.2 (parameters: $\beta=0.5$ and $\alpha=1.5$) with fixed block size $r=90$ and total sample size ranging form $40\cdot 90=3,600$ up to $100\cdot 90=9,000$. The target parameter is the runtime for calculating $B\in\{250,500,750,1,000\}$ bootstrap replicates of $\hat{\bm{\theta}}_{n}^{({\operatorname{mb}})}$ from (5.1) for ${\operatorname{mb}}\in\{{\operatorname{db}},{\operatorname{sb}},{\operatorname{cb}}(2),{\operatorname{cb}}(3)\}$, assessed by taking the median over $N=500$ repetitions each. The disjoint blocks method has been considered as a benchmark, whence we depict relative runtimes with respect to that method. We find that, as expected, the relative runtime is mostly depending on $n$, with only a moderate loss in performance for the circmax-method. Similar results were obtained for other estimators and models.
Figure 11: Relative median runtimes of different bootstrap algorithms for bootstrapping $\hat{\bm{\theta}}_{n}^{({\operatorname{mb}})}$ (relative to the runtime of the disjoint blocks bootstrap) for fixed blocksize $r=90$ as a function of the effective sample size and for different numbers of bootstrap replicates $B.$
Acknowledgements
Financial support by the German Research Foundation (DFG grant number 465665892) and by Ruhr University Research School (funded by Germany’s Excellence Initiative - DFG GSC 98/3) is gratefully acknowledged. Computational infrastructure and support were provided by the Centre for Information and Media Technology at Heinrich Heine University Düsseldorf. The authors are grateful to Johan Segers for fruitful discussions on the circular block maxima method, and to the participants of the Oberwolfach Workshop on “Mathematics, Statistics, and Geometry of Extreme Events in High Dimensions” for their valuable comments.

### v2

@@ F.4 Runtime comparison
The above heuristic has been confirmed by Monte Carlo simulations. Exemplary results are presented in Figure 24, which rely on simulated data from Model F.1 (parameters: $\beta=0.5$ and $\alpha=1.5$) with fixed block size $r=90$ and total sample size ranging form $40\cdot 90=3,600$ up to $100\cdot 90=9,000$. The target parameter is the runtime for calculating $B\in\{250,500,750,1,000\}$ bootstrap replicates of $\hat{\bm{\theta}}_{n}^{[{\operatorname{m}}]}$ from (A.1) for ${\operatorname{m}}\in\{{\operatorname{d}},{\operatorname{s}},{\operatorname{c}}\}$, assessed by taking the median over $N=500$ repetitions each. The disjoint blocks method has been considered as a benchmark, whence we depict relative runtimes with respect to that method. We find that, as expected, the relative runtime is mostly depending on $n$, with only a moderate loss in performance for the circmax-method. Similar results were obtained for other estimators and models.
Figure 24: Relative median runtimes of different bootstrap algorithms for bootstrapping $\hat{\bm{\theta}}_{n}^{[{\operatorname{m}}]}$ (relative to the runtime of the disjoint blocks bootstrap) for fixed blocksize $r=90$ as a function of the effective sample size and for different numbers of bootstrap replicates $B.$